# GET-Zero: Graph Embodiment Transformer for Zero-shot Embodiment Generalization

### v1

@@ 5 Experiments
  • ET. The Embodiment Transformer (ET) matches the architecture and token format of GET except without the graph encoding (§3.2) or the self modeling loss. This ET baseline most similarly matches the encoding method in Amorpheus [3], which uses no graph encoding. This means that the model has no explicit connectivity information about the joints.
  • ET+DFS. This baseline uses the same positional encoding scheme as MetaMorph [4] that linearizes the embodiment graph in depth-first search (DFS) ordering and then applies a learned positional encoding onto the linear sequence.
We also ablate different components of our method such as self-modeling loss (SL), spatial encoding (SE), parent-child encoding (PE). Results are summarized in Tab. 1.
Evaluation Metric. The task performance is measured by average rotational velocity along the yaw axis in degrees/second including standard deviation across five seeds. We compute results by rolling out the embodiment-aware policies in a physics simulator for a total of 42 minutes of execution time per embodiment per seed. Each category averages over 10 unseen embodiments (Appendix Fig. 6).
5.1 Key Results and Findings
@@
ET+PE+SE+SL (GET-Zero) $\bm{\pm}$0.24 $\bm{\pm}$0.58 $\bm{\pm}$0.29 $\bm{\pm}$0.54
Table 1: Simulation results (average rotational velocity deg/sec). RL experts are embodiment-specific policies. ET: Embodiment transformer, DFS: depth-first graph encoding, SL: self-modeling loss, SE: spatial graph encoding, PE: parent-child graph encoding. We observe GET-Zero consistently outperforms prior works [3, 4] through the use of graph encoding and self-modeling meta-task.
Table 1 summarizes GET-Zero’s performance on training embodiments (Training Graph) and on multiple categories of zero-shot embodiments: unseen graph variations (New Graph), unseen link length variations (New Geo) and both (New Graph & Geo). Note that link length variations (Fig. 4) were never seen during training.
GET-Zero matches the performance of embodiment-specific experts on training embodiments. Despite being trained to control 44 embodiments with graph changes, GET-Zero reaches 99% of the performance of the embodiment-specific RL experts with a single set of network weights (Tab. 1).
GET-Zero improves zero-shot capabilities to graph and geometry variations. Compared to the best performing baseline for each category, GET-Zero achieves a 16% improvement with zero-shot to new graph, an 10% improvement with zero-shot to link length variations, and a 20% improvement with zero-shot to both graph and link length variations (Tab. 1).
Self-modeling improves performance only with graph encoding present. We observe that adding self-modeling (SL) to the baseline (ET) consistently decreases zero-shot performance by an average of 4.6% across the three categories, but yields an average increase of 5.1% when added to the model with graph encoding (ET+PE+SE) (Table 1). This result intuitively makes sense as the ET baseline has no graph encoding which is needed to complete the forward kinematics task and thus self-modeling provides a poor training signal. In the presence of graph encoding, forward kinematics serves as a useful meta-task to improve cross-embodiment transfer.
Depth-first graph linearization lowers performance under graph variations. The DFS linearization (ET+DFS) serves as a simple, yet incomplete graph representation as the DFS ordering is not unique. Additionally, if a joint earlier in the linearization is removed, then all subsequent joints will shift down and receive a new positional encoding. When only link length variations are present, linearization improves zero-shot performance over no positional encoding (ET) by 11% (Tab. 1). However, when unseen graph changes are present, performance drops by 26%. This result intuitively makes sense as the linearized positional encoding can overfit to training embodiment graphs, but fails to generalize with graph changes due to the aforementioned issue with DFS. This motivates the need for the more complete graph representation used in GET-Zero.
Spatial and parent-child encoding help policy learning. Relative to ET with self-modeling (ET+SL), adding the parent-child graph encoding (ET+PE+SL) improves performance by 19% and adding the spatial graph encoding (ET+SE+SL) improves performance by 25% on average across zero-shot tasks. We note that adding parent-child encoding provides only a minor improvement when the spatial encoding is present, but that best performance is when both encodings are present.
Figure 5: Impact of training embodiments on zero-shot graph generalization.
Training diversity improves zero-shot. As shown in Fig. 5, increasing diversity of training embodiments improves zero shot generalization performance to unseen graph changes. GET-Zero also achieve a reasonable rotation performance even with only 10 training embodiments, and we observe that GET-Zero achieves more data efficient generalization performance than ET.
@@
Table 2: Real world evaluation. We evaluate sim-to-real capabilities of GET-Zero and the ET baseline on LEAP Hand [8] design variations. [Col 1] An AR tag tracks cube rotation. [Col 2] The unmodified LEAP Hand (seen during training). [Col 3-7] Zero-shot embodiments with unseen graph variations and/or link length extensions (orange). We report (across 3 minute trial): 1) average cube angular velocity in degrees/second 2) real average velocity as a % of sim average velocity for the same embodiment and policy (sim-to-real gap) and 3) number of times the hand dropped the cube.
5.2 Real-world Evaluation
We present results from a real-world evaluation of GET-Zero on unseen embodiments in Tab. 2. We observe GET-Zero outperforms the ET baseline for both the original LEAP Hand [8] as well as unseen graph and link length variations. Empirically, the baseline ET policy struggles to continuously rotate the cube, and we observe high finger precision is required for stable control. We observe a sim-to-real gap, but GET-Zero achieves zero-shot performance above simulation for the third embodiment. For another embodiment (second from right), a missing index finger causes the hand to drop the cube many times. For the right-most embodiment, a shortened index finger and thumb make it challenging for the hand to start the rotation cycle. Training with link length extensions could help improve zero-shot performance to New Graph+Geo. Please refer to the appendix for real-world evaluation methodology and the website for videos.
5.3 Limitations
While GET-Zero improves zero-shot performance, it’s unlikely that our model trained on LEAP Hand configurations would zero-shot transfer to a different robot hand. For example, we train with at most 16 joints, and expect poor performance for hands with more joints or fingers. Through compatible with the GET tokenization, our model does not explicitly encode joint limit ranges, motor strength, friction properties or finger shape, all of which vary with a new hand and are important in manipulation tasks. We also assume that the hand design space has enough variations to train on to expect reasonable generalization performance. Additionally, GET-Zero assumes a fixed set of embodiments, though future work could extend the model to iteratively improve a robot’s design.
6 Conclusion
We present GET-Zero, an embodiment-aware model architecture and training procedure that enables zero-shot control of new robot designs. Through a case study on an in-hand object rotation task, we demonstrate the ability of our model to control a wide range of hardware configurations of a multi-fingered robot hand under variations in embodiment graph and geometry. Our results in simulation and real demonstrate that the graph encoding and self-modeling features in GET-Zero improve cross-embodiment transfer. Our goal is for GET-Zero to serve as a useful, embodiment-aware method for the robotics community to share knowledge across similar robot designs.
Acknowledgments
We would like to thank Kenneth Shaw, Prof. Pathak’s lab at CMU, and the Prof. Liu’s Movement lab at Stanford for sharing the LEAP Hand hardware. We would also like to thank Huy Ha, Xiaomeng Xu, Mengda Xu and Haochen Shi for their helpful feedback and fruitful discussions. This work was supported in part by Sloan Fellowship and NSF #2132519. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.
A Appendix
We provide additional details on the procedural embodiment generation (§A.1), real-world evaluations (§A.2), a forward-kinematics task (§A.3), and the GET model and training parameters (§A.4).
A.1 Procedural Generation
In this section, we supplement the discussion in §4.1 with more details on the procedural generation of LEAP Hand embodiments. In Fig. 6 we show the diversity of zero-shot embodiments we use to evaluate GET-Zero spanning both graph and link length variations. We were able to train embodiment-specific RL experts for all 10 embodiments with graph variations (Fig. 6, row a). Doing so lets us verify that these hardware configurations are capable of completing the task with sufficient performance ($2\pi$ rotation completed in under 30 seconds), making them reasonable targets for zero-shot generalization.
We generate the zero-shot link length extension embodiments (Fig. 6 rows b and c) by introducing link extensions to the graph variation embodiments. In particular, for row b we randomly select 10 embodiments from our set of 44 training embodiments and introduce link extensions. For row c, we use the same 10 embodiments used to evaluate zero-shot graph generalization (row a). We open source all of our code for this work on our project website, including the scripts to procedurally generate embodiments as well as the generated URDF files and assets.
Figure 6: Zero-shot embodiments. We evaluate zero-shot transfer to (a) new graph variations, (b) link length extensions (orange), and (c) both link and graph variations. For each category we evaluate ten embodiments in simulation (pictured) and a subset to evaluate in real. Hands in the third row (link and graph variations) match those in the first row (connectivity variations), but with extensions applied to some links. We train on 44 hands with connectivity variations (not pictured).
A.2 Real-world Evaluation Methodology
We conduct real world evaluation across 3 minutes of execution per embodiment and sample our GET model at 20Hz. Every 20 seconds we pause execution, automatically reset the hand to a starting pose, and manually place the cube at the center of the palm. This matches the reset procedure used in the simulation evaluations and addresses the issue where the cube could get stuck in one challenging position for the entire evaluation.
Additionally, we attach an AR tag to the top of the cube and use the RGB camera in an Intel Realsense D415 camera to track the AR tag pose to provide the rotation metric. If the cube falls off the hand, the AR tag is no longer visible and we pause execution, increment the drop counter, reset the hand pose, and reset the cube.
A.3 Forward Kinematics Task
@@
To further understand the self-modeling component of GET, we train a GET model with only the forward kinematics prediction head and without the action head. The main objectives are to validate 1) that the model can leverage the embodiment graph to complete a downstream task and 2) that the learned graph encoding appropriately transfers to previously unseen embodiments. We observe that GET-Zero is able to achieve much higher precision predicting the per-joint positions because of the graph encoding mechanism, which provides joint graph connectivity information that is needed to accurately compute forward kinematics (Tab. 3).
@@
Table 3: Zero-shot forward kinematics performance. Measurements are a 5 seed average per-joint forward kinematics error in millimeters across 10 unseen embodiments with graph variations.
A.4 GET Model and Training Parameters
We include the parameter choices for the GET architecture, the behavior cloning training, and the dataset in Tab. 4. We refer to Shaw et al. [8] for the embodiment-specific model architecture and training parameters used for RL experts.
Parameter Value
Token dimension 512
Token dropout 0.1
Encoder layers 8
Attention heads 16
GET Model Attention dropout 0.1
Graph embedding initialization $\mathcal{N}(0,0.02)$
Feedforward dimension 1024
Head hidden dimension 16
Head activation elu
Action loss weight 1
Self-modeling loss weight 2
Optimizer Adam
Adam $\beta_{1}$, $\beta_{2}$ 0.9, 0.98
Behavior Cloning Epochs 7
LR schedule 15% Linear warmup with sqrt decay
LR max 1.5e-4
Grad clip 1
Training + sim eval time 8 hours on NVIDIA RTX A6000
Training embodiments 44
Dataset Train samples per embodiment 5e5
Zero-shot embodiments (Fig. 6) 30
Table 4: GET model, training and dataset parameters.

### v2

@@ V Experiments
  • ET. The Embodiment Transformer (ET) matches the architecture and token format of GET except without the graph encoding (§III-B) or the self modeling loss. This ET baseline most similarly matches the encoding method in Amorpheus [3], which uses no graph encoding. This means that the model has no explicit connectivity information about the joints.
  • ET+DFS. This baseline uses positional encoding from MetaMorph [4] that linearizes the embodiment graph in depth-first search (DFS) ordering and applies a learned linear positional encoding.
We also ablate different components of our method such as self-modeling loss (SL), spatial encoding (SE), parent-child encoding (PE). Results are summarized in Tab. I.
Evaluation Metric. The task performance is measured by average rotational velocity along the yaw axis in degrees/second including standard deviation across five seeds. We compute results by rolling out the embodiment-aware policies in a physics simulator for a total of 42 minutes of execution time per embodiment per seed. Each category averages over 10 unseen embodiments (see website for embodiment images).
Our expectation for zero-shot graph variation embodiments (given 44 training embodiments) is that joint sequences composing individual finger designs will have been seen before, but that 1) the finger position within the hand will be unseen (three positions corresponding to the three main fingers) and/or that 2) the combinations of finger designs across the four fingers will be unseen. We were able to train embodiment-specific RL experts for all 10 embodiments with graph variations with sufficient performance ($2\pi$ rotation within 30 seconds), validating that they are reasonable targets for zero-shot generalization.
TABLE I: Simulation results (average rotational velocity deg/sec). RL experts are embodiment-specific policies. ET: Embodiment transformer, DFS: depth-first graph encoding,
SL: self-modeling loss, SE: spatial graph encoding,
PE: parent-child graph encoding
@@
ET+PE+SE+SL $\bm{\pm}$0.24 $\bm{\pm}$0.58 $\bm{\pm}$0.29 $\bm{\pm}$0.54
TABLE II: GET-Zero sim-to-real evaluation. [Col 1] An AR tag tracks cube rotation. [Col 2] The unmodified LEAP Hand (seen during training). [Col 3-7] Zero-shot embodiments with unseen graph variations and/or link length extensions (orange). We report (across 3 minute trial): 1) average cube angular velocity in degrees/second 2) real average velocity as a % of sim average velocity for the same embodiment and policy (sim-to-real gap) and 3) number of times the cube fell off the hand.
@@
V-A Key Results and Findings
Table I shows GET-Zero’s performance on training embodiments (Training Graph) and on multiple types of zero-shot embodiments: unseen graph variations (New Graph), unseen link length variations (New Geo) and both (New Graph & Geo). Link length extensions were never seen during training.
GET-Zero matches the performance of embodiment-specific experts on training embodiments. Despite controlling 44 embodiments with graph changes using a single set of network weights, GET-Zero reaches 99% of the performance of the embodiment-specific RL experts (Tab. I).
GET-Zero improves zero-shot capabilities to graph and geometry variations. Compared to the best performing baseline for each category, GET-Zero achieves a 16% improvement with zero-shot to new graph, an 10% improvement with zero-shot to link length variations, and a 20% improvement with zero-shot to both graph and link length variations (Tab. I).
Self-modeling improves performance only with graph encoding present. We observe that adding self-modeling (SL) to the baseline (ET) consistently decreases zero-shot performance by an average of 4.6% across the three categories, but yields an average increase of 5.1% when added to the model with graph encoding (ET+PE+SE) (Table I). This result intuitively makes sense as the ET baseline has no graph encoding which is needed to complete the forward kinematics task and thus self-modeling provides a poor training signal. In the presence of graph encoding, forward kinematics serves as a useful loss to improve cross-embodiment transfer.
Depth-first graph linearization lowers performance under graph variations. The DFS linearization (ET+DFS) serves as a simple, yet incomplete graph representation as the DFS ordering is not unique. Additionally, if a joint earlier in the linearization is removed, all subsequent joints will shift down and receive a new positional encoding. When only link length variations are present, linearization improves zero-shot performance over no positional encoding (ET) by 11% (Tab. I). However, with unseen graph variations, performance drops by 26%. This result intuitively makes sense as the linearized positional encoding overfit to training embodiment graphs, but fail to generalize to unseen graphs due to the aforementioned issue with DFS. This motivates the need for the more complete graph representation used in GET-Zero.
Spatial and parent-child encoding help policy learning. Relative to ET with self-modeling (ET+SL), adding the parent-child graph encoding (ET+PE+SL) improves performance by 19% and adding the spatial graph encoding (ET+SE+SL) improves performance by 25% on average across zero-shot tasks. Adding the parent-child bias when the spatial bias is already present provides no statistically significant improvement, indicating that our directed encoding is not critical for performance, perhaps as parent-child relationships can be inferred from correlations in joint states.
Figure 5: Impact of training embodiments on zero-shot graph generalization. We observe that even with fewer training embodiments, GET-Zero achieves reasonable in-hand rotation performance (5 seeds).
Zero-shot with fewer training embodiments. To validate that zero-shot performance isn’t solely due to a large number of training embodiments that might be similar to the zero-shot embodiments, we evaluate our method on smaller subsets of training embodiments as shown in Fig. 5. GET-Zero achieves reasonable rotation performance even with only 10 training embodiments, indicating that zero-shot performance works when demonstration data is only available for fewer embodiments. The 44 training embodiments for results in Tab. I is 19% of the 236 total graph variation embodiments generated, though not all 236 embodiments are equally capable of completing the rotation task.
@@ V-A Key Results and Findings
Graph encoding improves forward kinematics prediction. We train a GET model with only the forward kinematics head (no action head) to validate self-modeling capabilities for unseen embodiments. In Tab. III we observe that GET-Zero achieves much higher precision predicting per-joint positions due to the graph encoding mechanism.
V-B Real-world Evaluation
We present results from a real-world evaluation of GET-Zero on unseen embodiments in Tab. II following the same evaluation methods as for simulation. We observe GET-Zero outperforms the ET baseline for both the original LEAP Hand [13] as well as unseen graph and link length variations. Empirically, the ET policy struggles to continuously rotate the cube, and we observe high finger precision is required for stable control. We observe a sim-to-real gap, but GET-Zero achieves zero-shot performance above simulation for the third embodiment. For another embodiment (second from right), a missing index finger causes the hand to drop the cube many times. For the right-most embodiment, a shortened index finger and thumb make it challenging for the hand to start the rotation cycle.
TABLE III: Average per-joint forward kinematics error in millimeters
across 10 zero-shot graph variations (5 seeds)
@@ V-B Real-world Evaluation
V-C Limitations
It’s unlikely that GET-Zero trained on LEAP hands would zero-shot transfer to a new robot hand model. For example, we train with at most 16 joints and do not encode joint limit ranges, motor strength, friction properties or finger shape, all of which vary with a new hand and are important in manipulation tasks. Future work can extend GET-Zero to learn to iteratively improve a robot’s design or explore if GET-Zero supports multi-task settings such as locomotion and manipulation with a single network.
VI Conclusion
We present GET-Zero, an embodiment-aware model architecture and training procedure that enables zero-shot control of new robot designs. Through a case study on an in-hand object rotation task, we demonstrate the ability of our model to control a wide range of hardware configurations of a multi-fingered hand under variations in embodiment graph and geometry. Our results in simulation and real show that the graph encoding and self-modeling features in GET-Zero improve cross-embodiment transfer. We hope that GET-Zero serves as a useful, method for the robotics community to share knowledge across similar robot designs.
ACKNOWLEDGMENT
We thank Kenneth Shaw, Prof. Pathak’s lab at CMU, and the Prof. Liu’s Movement lab at Stanford for sharing LEAP Hand hardware. We also thank Huy Ha, Xiaomeng Xu, Mengda Xu and Haochen Shi for their helpful feedback and fruitful discussions. This work was supported in part by Sloan Fellowship and NSF #2132519. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.